OpenStudy (anonymous):

if sec theda=7, find cos theda

Vocaloid (vocaloid):

hint: cos(theta) = 1/sec(theta)

hero (hero):

First time I ever saw anyone spell theta with a d

OpenStudy (anonymous):

oops i meant theta @Hero

hero (hero):

@cali16, what is the relationship between sec theta and cos theta?

OpenStudy (anonymous):

they both use the hypothenuse

hero (hero):

I was thinking more in terms of the equations that represent both trig functions.

OpenStudy (anonymous):

they are opposites

OpenStudy (anonymous):

so would it be 1/7

hero (hero):

They're INVERSES of each other, not OPPOSITES. But yes, cos theta = 1/7 if sec theta = 7.

OpenStudy (anonymous):

okay thanks